OpenStudy (anonymous):

Divide the following polynomial and then place the answer in the proper location on the grid. Write your answer in order of decreasing powers of x. (24x 5 + 8x 4 - 16x 3) ÷ 8x 2

OpenStudy (campbell_st):

easiest way is to write it as 3 fractions \[\frac{24x^5}{8x^3}+\frac{8x^4}{8x^3} - \frac{16x^3}{8x^3}\] simplify each term when dividing like terms the powers are subtracted 1st term becomes \[\frac{24x^5}{8x^3} = 3x^2\] I'll let you do the rest

OpenStudy (anonymous):

i have no idea

OpenStudy (campbell_st):

oops... got the denominator powers wrong \[\frac{24x^5}{8x^2} + \frac{8x^4}{8x^2} - \frac{16x^3}{8x^2}\]

OpenStudy (campbell_st):

1st term is \[\frac{24x^5}{8x^2} = 3x^{5-2} = 3x^3\]

OpenStudy (anonymous):

?

OpenStudy (campbell_st):

the next term is and remember 8/8 = 1 \[\frac{8x^4}{8x^2} = x^{4-2} = x^2\]